Chlorine trifluoride (ClF3) is used as a fluorinating agent. It may be used as an igniter and propellant in rockets and also used in nuclear fuel processing. Breathing Chlorine Trifluoride can irritate the lungs causing coughing and/or shortness of breath. Higher exposures can cause a build-up of fluid in the lungs (pulmonary edema), a medical emergency, with severe shortness of breath. Contact with the liquified gas can cause frostbite.


Using the given table below, formulate your own kind of problem; through providing hypothetical value for individual reactions, and also provide answer using the steps in getting the ΔH° of the reaction. ONLY condition is the reaction must be exothermic. 


ClF3(g) + O2 (g) → ½ Cl2O (g) + 3/2OF2 (g) ΔH° =

2ClF(g) + O2 (g) → Cl2O (g) + OF2 (g) ΔH° =

2OF2 (g) → O2 (g) + 2F2 (g) ΔH° = ClF (g) + F2 (g) → ClF3 (g) ΔH° = ?


1
Expert's answer
2021-05-17T04:11:35-0400

ClF3(g) + O2 (g) → ½ Cl2O (g) + 3/2OF2 (g) ΔH° = 170.7 kJ/mol

2ClF(g) + O2 (g) → Cl2O (g) + OF2 (g) ΔH° = 83.7 kJ/mol

2OF2 (g) → O2 (g) + 2F2 (g) ΔH° = ClF (g) + F2 (g) → ClF3 (g) ΔH° = 43.4 kJ/mol